points 60, $1635 120, $1770 I got a slope of 2.25 My intercept is 1,500. Not sure if this is right. Let x represent packages and y cost. What is the cost to send 200? What would my equation look like? Thanks
Answer by stanbon(75887)

points ( 60, $1635) (120, $1770_
I got a slope of 2.25
That is correct:
---------------------
My intercept is 1,500.
Substitute into y = mx+b to solve for "b"
1635 = 2.25*60 + b
b = 1500
You are correct:
---------------------
Not sure if this is right. Let x represent packages and y cost. What is the cost to send 200? What would my equation look like?
EQUATION:
y = 2.25x + 1500
To send 200 you get y = 2.25*200 + 1500
y = 450 +1500
y = 1950
===============
Cheers,
Stan H.